how would i do the following problem 2/3+1/2 divided by 3/4

Answer by oberobic(2304)

(2/3+1/2) / (3/4) = (4/6 + 3/6) / (3/4) = (7/6)/(3/4) = (7/6)*(4/3) = 28/18 = 14/9